### v1

# ANUBHUTI: A Comprehensive Corpus for Sentiment Analysis in Bangla Regional Languages

Abstract
Sentiment analysis for regional dialects of Bangla remains an underexplored area due to linguistic diversity and limited annotated data. This paper introduces ANUBHUTI, a comprehensive dataset consisting of 2,000 sentences manually translated from standard Bangla into four major regional dialects—Mymensingh, Noakhali, Sylhet, and Chittagong. The dataset predominantly features political and religious content, reflecting the contemporary socio-political landscape of Bangladesh, alongside neutral texts to maintain balance. Each sentence is annotated using a dual annotation scheme: (i) multiclass thematic labeling categorizes sentences as Political, Religious, or Neutral, and (ii) multilabel emotion annotation assigns one or more emotions from Anger, Contempt, Disgust, Enjoyment, Fear, Sadness, and Surprise. Expert native translators conducted the translation and annotation, with quality assurance performed via Cohen’s Kappa inter-annotator agreement, achieving strong consistency across dialects. The dataset was further refined through systematic checks for missing data, anomalies, and inconsistencies. ANUBHUTI fills a critical gap in resources for sentiment analysis in low-resource Bangla dialects, enabling more accurate and context-aware natural language processing.
Keywords Regional Dialects  $\cdot$ Sentiment Analysis  $\cdot$ Low Resource Language  $\cdot$ Bangla BERT
Specification Table
Subject Computer Science Applications
Specific subject area Sentiment Analysis on Bangla Regional text data.
Type of data Table (CSV)
Data collection ANUBHUTI comprises approximately 500 sentences per region. The raw sentences were collected from a publicly available Bangla Sentiment Analysis corpus and further manual transaltion was done via human annotators. The texts primarily cover religious and political content.
Data source location MONOVAB
Data accessibility Repository Name: Mendeley Data (10.17632/gbkszkt8z3.2)
Related research article None
1 Value of the data
  • ANUBHUTI is a pioneering corpus developed for sentiment analysis in Bangla regional dialects, addressing a critical gap in low-resource NLP research for languages such as Sylhet, Chittagong, Noakhali, and Mymensingh.
  • The dataset is thoroughly annotated by expert linguists specializing in regional Bangla dialects, ensuring high-quality sentiment analysis dataset.
  • The corpus provides a foundation for practical applications such as dialect-aware chatbots, mental health analysis, and social media monitoring, contributing to inclusive and regionally adaptive AI technologies.
  • Provided in a structured CSV format, the dataset is compatible with various NLP pipelines, making it a valuable resource for future academic and industry research in Bangla language processing.
2 Related Works
Sentiment and emotion analysis in morphologically rich and low-resource languages has witnessed considerable growth in recent years. However, the majority of progress has been concentrated around high-resource languages or standardized variants, leaving dialectal and regionally diverse linguistic forms underexplored. This literature review synthesizes recent studies across Bangla, Arabic, and Hindi, with a particular emphasis on dialect-aware sentiment classification, resource development, and methodological innovation.
2.1 Bangla Sentiment and Emotion Analysis
Bangla, as a low-resource yet widely spoken language, has increasingly become the focus of sentiment and emotion analysis efforts. Bhowmik et al. (2021)[1] proposed a supervised learning framework combining a categorical weighted lexicon (LDD) with a rule-based sentiment scoring method called Bangla Text Sentiment Score (BTSC). Applied to cricket and restaurant review datasets, their SVM model achieved a classification accuracy of 82.21% using BiGram features. However, the study underscored the shallow progress of Bangla sentiment analysis due to a lack of computational resources and annotated corpora.
To address the need for large-scale data, Kabir et al. (2023)[2] introduced BanglaBook, a dataset of over 158,000 Bangla book reviews categorized into positive, neutral, and negative sentiments. The authors evaluated multiple traditional and deep learning models, including fine-tuned Bangla-BERT, which achieved a weighted F1-score of 0.933. The work emphasized the shortage of domain-specific annotated datasets and advanced pre-trained models tailored for Bangla.
In a machine-learning-focused study, Paul et al. (2025)[3] analyzed Bangla social media comments using ML classifiers, BiLSTM, and LIME for explainability. Leveraging the EmoNoBa dataset with six emotion classes, their Decision Tree model boosted with AdaBoost achieved an F1-score of 0.786. The study addressed the concern that social media–based datasets might not reflect the full spectrum of Bangla linguistic usage.
In another recent study, Mahmud et al.(2024)[4] presented a benchmark dataset for sentiment analysis of Bangla cricket-related social media content, focusing on low-resource and noisy textual environments. With a Cohen’s Kappa score of 0.971 indicating strong annotation reliability, the dataset supported evaluations using deep learning models. The BiLSTM model achieved 95.2% accuracy, outperforming both LSTM and traditional RNN models. This work addressed the scarcity of annotated Bangla datasets in the sports domain.
Islam et al. (2023)[5] further advanced resource development by releasing SentiGOLD, a 70,000-sample Bangla dataset spanning over 30 domains such as news, blogs, and social media. Featuring a nuanced five-class sentiment scale, SentiGOLD achieved an Inter-Annotator Agreement of 0.88. Experiments showed that BanglaBERT performed competitively both in in-domain classification and zero-shot cross-dataset evaluations, outperforming prior datasets like SentNoB. The authors addressed issues of scale, bias, and annotation quality prevalent in earlier Bangla resources.
Beyond sentiment, Faisal et al. (2024)[6] explored emotion classification in their Bengali & Banglish dataset, which contains over 80,000 samples annotated across six emotional states. By combining data from EmoNoBa, UBMEC, and newly scraped social content, the authors captured both standard Bangla and Romanized Bangla (Banglish). Transformer models like BanglaBERT achieved an F1-score of 71.3% for Bangla and 64.6% for Banglish, revealing the linguistic complexity introduced by code-switching.
In the e-commerce domain, Haque et al. (2018)[7] analyzed sentiment in a large-scale Amazon product review dataset of over 48,000 entries across categories like electronics and musical instruments. Using active learning for data labeling and applying SVM with TF-IDF and n-gram features, the authors achieved over 90% classification accuracy. However, they noted the challenge of acquiring gold-standard annotated corpora from commercial sources.
2.2 Sentiment Analysis in Arabic Dialects
A foundational contribution in dialectal Arabic sentiment analysis is the work by Baly et al. (2017)[8], who introduced the Multi-Dialect Arabic Sentiment Twitter Dataset (MD-ArSenTD), containing tweets from 12 Arab countries with a focus on Egyptian and UAE dialects. The study compared Support Vector Machines (SVM) with engineered features against deep learning models like Long Short-Term Memory (LSTM). Results showed that LSTM models, when supplemented with dialect-specific word embeddings, outperformed traditional methods. Egyptian tweets, being highly dialectal, benefited more from morphological and localized lexical features, while UAE tweets were more aligned with Modern Standard Arabic (MSA). The work was the first to conduct sentiment analysis on UAE tweets and applied a novel five-point sentiment scale for Egyptian tweets.
Similarly, Mdhaffar et al. (2017)[9] tackled sentiment analysis for the Tunisian dialect, creating the first publicly available annotated sentiment corpus for Tunisian Arabic (TSAC), consisting of 17,000 Facebook comments. They experimented with classifiers including MLP, Naive Bayes, and SVM, and demonstrated that models trained on dialect-specific data yielded significantly better performance than those trained on other dialects or MSA. The study emphasized the need for tailored linguistic resources and highlighted the scarcity of NLP tools for North African Arabic dialects.
Table 1: Summary of Sentiment Analysis Study Across Languages and Datasets
Author Dataset Approaches Accuracy Baly et al.[8] MD-ArSenTD SVM, LSTM Egypt: SVM: 60.6%, LSTM: 70.0%
UAE: SVM: 51.1%, LSTM: 63.7%
Medhaffar et al.[9] TSAC SVM, Naive Bayes, MLP MLP: 78%, SVM: 77%, BNB: 58% Faria et al.[10] Vashantor mT5, BanglaT5, mBERT, Bangla-bert-base Bangla-bert: 85.86%, mBERT: 84.36% Pandey et al.[11] Hindi Review Dataset HSWN, Rule Handling HSWN-based: 80% Bhowmik et al.[1] Bangla ABSA UniGram, BiGram UniGram: 77.91%–78.69%
BiGram: 80.58%–82.21%
Kabir et al.[2] BanglaBook SVM, LSTM, Bangla-BERT Bangla-BERT (F1): 93.31% Mahmud et al.[4] Bangla Cricket Sentiment ML, DL (LR, RF, SVM, RNNs) RNN: 52.39%, LSTM: 89.72%, BiLSTM: 95.12% Islam et al.[5] SentiGOLD LSTM, BanglaBERT Macro F1: 0.62 (5-class), 0.61 (cross-set) Faisal et al.[6] Bangla + Banglish Machine Learning, Deep Learning Bangla: 71.12%, Banglish: 64.31% Haque et al.[7] Amazon Reviews NB, SVM, RF, etc. Music: 94.02%, Phones: 93.57%, Electronics: 93.52% Paul et al.[3] EmoNoBa n-gram, Machine Learning models: Decision Tree, SVM, Random Forest, KNN Decision Tree: 78.60%
2.3 Sentiment Analysis in Hindi
In the context of Hindi, Pandey and Govilkar (2015)[11] proposed an unsupervised sentiment analysis framework leveraging HindiSentiWordNet (HSWN). While effective for basic polarity detection, the HSWN lexicon was limited in coverage—particularly in handling adjectives and adverbs—and struggled with context-dependent word interpretation. The paper noted that without sufficient lexical expansion and automated context handling, accurate sentiment extraction in Hindi remains challenging.
2.4 Dialect Based Studies
A major enabling resource for dialectal Bangla NLP is Vashantor, proposed by Faria et al. (2023)[10]. This benchmark dataset facilitates the automated translation of regional Bangla dialects into standard Bangla. Covering a wide geographic and linguistic span—including Sylhet, Noakhali, Barishal, Chittagong, and Mymensingh dialects—Vashantor is a pivotal contribution toward dialect normalization and low-resource translation. Although not focused directly on sentiment analysis, the dataset provides foundational infrastructure to build dialect-aware NLP models and enhance downstream tasks such as emotion or sentiment classification in regional varieties of Bangla. Similar to Vashantor dataset, Sultana et al. (2025)[12] and Paul et al. (2025)[13] proposed their own version of Dialect Translation dataset. In a recent study of Regional Dialect, Paul et al. (2024)[14] explored regional dialect classification using data from the VASANTOR dataset, covering Bangla dialects like Mymensingh, Chittagong, Barishal, Noakhali, and Sylhet. The study compared BanglaBERT, GPT-3.5 Turbo, and Gemini 1.5 Pro, finding that BanglaBERT achieved the highest classification accuracy of 88.74%. This work underlines the complexity of intra-language variation and the need for dialect-sensitive models.
Table 1 summarizes the literature reviews of our study. Despite significant advancements in sentiment and emotion analysis for standard Bangla, Hindi, and Arabic dialects, a notable research gap persists in the area of regional dialect sentiment analysis, particularly for low-resource languages like Bangla. Existing studies either rely heavily on standard language forms or lack the granularity to capture regional linguistic nuances. While efforts like Vashantor and regional classification models have laid the groundwork, sentiment-specific datasets and models for Bangla dialects remain virtually nonexistent. This underscores the urgent need to develop sentiment analysis systems that can understand and process non-standard dialects, which are prevalent in real-world communication but poorly represented in current NLP infrastructure. In response to this gap, our research aims to advance dialect-aware sentiment analysis by leveraging modern transformer architectures, domain adaptation, and interpretability techniques—thus contributing to a more inclusive and linguistically representative natural language processing landscape.
3 Data Description
Sentiment analysis, also known as opinion mining, is a fundamental task in Natural Language Processing (NLP) that involves identifying and categorizing emotions or opinions expressed in text. It plays a vital role in a wide range of applications, including social media analysis, political forecasting, and consumer behavior research, by assessing the polarity of a given textual input.
Although sentiment analysis has achieved substantial progress in high-resource languages like English, its development in low-resource languages such as Bangla is still in the early stages. Most existing works in Bangla sentiment analysis focus primarily on formal or standard language forms, overlooking the rich diversity of dialectal variations that exist across different regions of Bangladesh. However, dialectal expressions are often more emotionally charged, contextually grounded, and reflective of grassroots-level sentiment. Addressing this gap, our work introduces a new perspective by focusing on Bangla regional dialects for sentiment analysis. This approach enables a more inclusive and granular understanding of public opinion, especially in socio-political and religious contexts that dominate online discourse in Bangladesh.
To support this goal, we developed a dialect-sensitive sentiment dataset named ANUBHUTI, which is tailored to capture the linguistic and emotional nuances present in regional Bangla. Figure 1 illustrates the development pipeline of the ANUBHUTI dataset, highlighting the key stages of data sourcing, filtering, preprocessing, and annotation.
Figure 1: A Systematic Pipeline for the development of ANUBHUTI Dataset
3.1 Data Collection
The dataset used in this study, titled ANUBHUTI, was developed using raw data sourced from the publicly available MONOVAB dataset. MONOVAB comprises a diverse range of user-generated Bangla content, making it a valuable resource for mining sentiment-rich expressions in regional dialects.
For the initial phase of this research, a total of 500 raw data samples were collected to support preliminary sentiment analysis. The primary focus was on political and religious discourse, as these domains tend to evoke strong emotional reactions and diverse public opinions. A subset of neutral content was also included to provide a baseline for comparative sentiment evaluation. The choice of political and religious content was motivated by the current socio-political climate in Bangladesh, where heightened public discourse around governance, ideological tensions, and religious sentiments is prominently visible across digital platforms. These topics often reflect deeper emotional undercurrents and are highly relevant for training sentiment classification models that aim to capture context-sensitive regional expressions.
3.2 Data Translation
To construct a regionally diverse sentiment dataset, we selected four prominent dialect regions of Bangladesh: Mymensingh, Noakhali, Sylhet, and Chittagong. These regions were chosen based on their distinct linguistic features and strong presence in local digital discourse. To ensure balanced representation, we translated 500 sentences into the corresponding dialect for each region, resulting in a total of 2,000 dialectal samples.
Given the complexity and variability of regional expressions, two native dialect experts were hired for each region to carry out the translation task. These experts possess both linguistic familiarity and contextual understanding of local expressions, allowing them to preserve the semantic integrity and sentiment polarity of each sentence during translation. The collaborative process between paired experts ensured consistency, resolved ambiguities, and maintained dialectal authenticity. This translation step was crucial for adapting sentiment-labeled content into natural, expressive dialectal forms that reflect real-world language use. The translated data serves as a foundation for training sentiment analysis models that can operate effectively on informal and region-specific Bangla text.
3.3 Data Translators Identity
To ensure the linguistic and contextual accuracy of dialectal translations, we recruited eight native speakers with proven expertise in their respective regional dialects—two for each of the selected regions: Mymensingh, Noakhali, Sylhet, and Chittagong. All translators were either language professionals, postgraduate students in linguistics, or individuals with prior experience in language documentation or translation work. Their strong familiarity with regional speech patterns, idiomatic expressions, and sentiment articulation made them highly suitable for the task. A thorough screening process was conducted to assess their dialectal proficiency and ensure consistency in translation quality. Detailed information about the translators is presented in Table 2.
Table 2: Background and Expertise of Dialect Translators
Region ID Background Expertise
Mymensingh T1, T2 Under GraduateStudent Native Speaker, Translation Experience
Noakhali T3, T4 GraduateStudent, Journalist Native Speaker
Sylhet T5, T6 Researcher Native Speaker
Chittagong T7, T8 UnderGraduate Native Speaker
3.4 Cohen’s Kappa for Translation Agreement
To ensure the quality and reliability of dialectal translations in the ANUBHUTI dataset, it was important to evaluate the inter-annotator agreement between the two translators assigned to each region. Given the subjective nature of language translation, particularly in dialects with no standardized written form, a robust agreement metric was essential.
For this purpose, Cohen’s Kappa ($\kappa$) was selected as a statistical measure to quantify the level of agreement between two raters beyond chance. Unlike simple percentage agreement, Cohen’s Kappa accounts for the probability of agreement occurring by chance, making it a more reliable metric for evaluating subjective tasks such as translation validation. A high $\kappa$ score indicates strong consistency and shared understanding between the translators, which is critical for maintaining sentiment and semantic fidelity in the dataset. As shown in Table 3, the scores for all four regions—Mymensingh, Noakhali, Sylhet, and Chittagong—fall within the range of 0.76 to 0.84, indicating substantial to almost perfect agreement.
The computation was based on a comparison between the annotations provided by the two translators per region, using the standard Cohen’s Kappa formula. For each sentence, agreement was recorded if both annotators provided semantically and sentimentally aligned translations. The observed agreement ($p_{o}$) and the expected agreement by chance ($p_{e}$) were computed from the annotation distribution, and Cohen’s Kappa ($\kappa$) was derived using:
$\kappa=\frac{p_{o}-p_{e}}{1-p_{e}}$
Where:
  • $p_{o}$ is the observed agreement between annotators,
  • $p_{e}$ is the expected agreement by chance.
Table 3: Cohen’s Kappa Scores for Translation Agreement by Region
Region Kappa Score ($\kappa$)
Mymensingh 0.84
Noakhali 0.78
Sylhet 0.81
Chittagong 0.76
These region-specific scores confirm a high degree of consistency and mutual understanding between translators, reinforcing the quality and reliability of the ANUBHUTI dataset’s dialectal representations.
3.5 Data Inconsistency Checking
In the context of this research on sentiment analysis using regional dialects of Bangla, ensuring consistency and quality across all translated data is a crucial step. Since the original dataset was composed in standard Bangla and manually translated into four regional dialects—Mymensingh, Noakhali, Sylhet, and Chittagong—the risk of inconsistency was inherently high due to linguistic variability, subjective translation choices, and human error. To address these challenges, a systematic and multi-stage data inconsistency checking framework was developed. This framework consists of four primary stages:
3.5.1 Re-evaluation
In this phase, all translated sentences were manually reviewed and cross-checked against the original standard Bangla texts. The goal was to ensure semantic alignment and sentiment preservation. Reviewers focused on maintaining idiomatic correctness, dialect-specific vocabulary, and the emotional tone of each sentence. This process helped guarantee that the translated dialectal texts accurately reflected the intent and polarity of the original content.
3.5.2 Missing Data Detection
To ensure the completeness and reliability of the translated dataset, a dual-phase approach was adopted for identifying missing or incomplete entries. This involved both manual inspection and automated detection using Python scripting.
The manual review process was conducted to check for:
  • Entirely absent translations,
  • Partially completed sentences,
  • Omission of key sentiment-bearing words or phrases.
Each translated entry was validated against its original Bangla source to ensure semantic completeness and fidelity of sentiment expression, thereby reducing the risk of information loss during translation.
To complement this, a Python-based script was developed to automatically identify missing values within the dataset. The script reads the dataset using the pandas library, checks each column for NaN or null values, and summarizes the missing data count. If any missing entries are detected, the script provides a detailed report for further review. The core logic of this automated detection system is outlined in Algorithm 1.
Algorithm 1 Check for Missing Data in CSV File
1:procedure CheckMissingData(filepath)
2:     Import the pandas library
3:     df $\leftarrow$ pandas.read_csv(filepath)
4:     missing_values $\leftarrow$ df.isnull().sum()
5:     total_missing $\leftarrow$ df.isnull().sum().sum()
6:     if total_missing > 0 then
7:         Print “Missing values found:”
8:         Print missing_values
9:     else
10:         Print “No missing values found.”
11:     end if
12:end procedure
3.5.3 Anomaly Detection
To ensure the quality and dialectal integrity of the translated dataset, an anomaly detection phase was implemented. This phase aimed to identify irregularities that deviated from the expected linguistic and contextual norms of each target dialect. Anomalies were broadly categorized into two types: labeling inconsistencies and textual anomalies.
Manual review was used to detect issues such as:
  • Sentences that mixed multiple dialects or included standard Bangla unintentionally,
  • Translations that were contextually inappropriate or semantically misaligned with the original sentence,
  • Use of grammatical constructions that did not conform to the syntax of the target dialect.
In addition to manual checking, a semi-automated Python-based anomaly detection script was developed to systematically flag problematic entries. The script operated according to the following step, where the system is evaluated sentence formatting to identify unusual patterns such as excessive punctuation (e.g., “!!!!”, “….”), non-Bangla characters or scripts, and other spam-like structures. Such anomalies were considered indicators of noise or stylistic inconsistency that could impact downstream analysis.
All automatically flagged entries were reviewed manually to confirm their anomalous status, and necessary corrections were applied to preserve semantic, emotional, and dialectal integrity. The structured logic of this anomaly detection procedure is presented in Algorithm 2.
Algorithm 2 Detect Anomalies in Bangla Sentiment Dataset
1:procedure DetectAnomalies(Dataset)
2:     Import pandas, re, sentence-transformers, sklearn, fuzzywuzzy
3:     df $\leftarrow$ Load the CSV file
4:
5:     for each text in df[‘text’] do
6:         if Contains repeated chars (e.g., “!!!!” or “….”) or non-Bangla scripts then
7:              Flag as Content Anomaly: Suspect formatting or language
8:         end if
9:     end for
10:end procedure
3.5.4 Spelling Verification
Due to the lack of standardized orthography for many regional Bangla dialects, spelling verification was performed manually with the aid of native speakers. The goal was to ensure:
  • Consistency in commonly accepted dialectal spellings,
  • Phonetic alignment with regional pronunciation,
  • Elimination of ambiguities caused by irregular spelling variants.
Given the informal nature of written regional dialects, this step was essential in preserving clarity, authenticity, and interpretability in the translated texts.
The manual inconsistency checking process played a foundational role in enhancing the reliability and validity of the ANUBHUTI dataset. By ensuring translation accuracy, completeness, contextual consistency, and dialect-specific spelling correctness, the dataset was rendered suitable for training robust and dialect-aware sentiment analysis models. This quality control phase ensures that the models can effectively interpret and classify emotional content in diverse forms of Bangla, particularly those originating from underrepresented regional communities.
3.6 Data Annotation
To prepare the ANUBHUTI dataset for supervised sentiment analysis, a two-level annotation process was carried out. The annotation involved manual labeling of each sentence for both its thematic class (multiclass) and emotional sentiment (multilabel). These annotations enable the development and evaluation of models capable of understanding complex emotional expression and contextual bias within Bangla regional dialects. The decision to implement both multiclass and multilabel annotation schemes was motivated by the need to capture the full complexity of sentiment expression in regional Bangla dialects. The multiclass annotation provides a high-level thematic categorization—distinguishing whether a sentence is political, religious, or neutral—allowing for context-aware analysis of sentiment trends. Meanwhile, the multilabel emotion tagging captures the diverse and overlapping emotional states conveyed in informal dialectal speech.
3.6.1 Multiclass Annotation
Each sentence in the dataset was annotated with one of three mutually exclusive thematic labels: Political, Religious, or Neutral. This multiclass classification captures the overarching context or topic of the sentence, and is crucial for understanding domain-specific sentiment patterns.
  • Political: Includes references to governance, elections, political figures, movements, or actions.
  • Religious: Encompasses content related to faith, belief systems, religious practices, and figures.
  • Neutral: Covers general content without explicit political or religious context.
Each sentence was assigned exactly one of these classes, ensuring that thematic context is represented clearly and non-overlappingly. This classification supports topic-aware sentiment analysis and can be used to analyze how sentiment varies across different socio-cultural domains.
3.6.2 Multilabel Annotation
In addition to contextual classification, each sentence was manually annotated for emotional sentiment using a multilabel scheme. The following seven core emotions were used as annotation labels:
  • Anger: Assigned to sentences that express frustration, outrage, or hostility, often directed at political figures, events, or institutions. For example, accusatory or confrontational statements about government actions or religious intolerance.
  • Contempt: Applied to sarcastic, belittling, or demeaning expressions — particularly those mocking political leaders, religious opponents, or ideologies. Contemptuous tone is often implicit and requires contextual understanding.
  • Disgust: Used for sentences reflecting moral revulsion, societal disappointment, or emotional rejection. This often includes strong criticisms of corruption, injustice, or offensive actions within political or religious contexts.
  • Enjoyment: Assigned to sentences expressing positive emotion, such as approval, support, or satisfaction. In political or religious contexts, this might include praise for actions, victories, or celebrations of community events.
  • Fear: Used for sentences that indicate anxiety, insecurity, or anticipation of negative consequences. This often includes references to political instability, religious violence, or threats to personal or communal safety.
  • Sadness: Assigned when the sentence reflects grief, loss, or sorrow — such as mentions of deaths, tragedies, or despair related to social or political events.
  • Surprise: Applied to sentences indicating unexpected outcomes or shock — whether positive or negative. This may involve disbelief at political scandals, sudden policy changes, or religious controversies.
A sentence could exhibit one, multiple, or even no emotion labels depending on its emotional tone. This multilabel approach captures the nuance and complexity of real-life emotional expression, especially within dialectal text where mixed sentiment is common. For example, a politically charged sentence might simultaneously express both Anger and Contempt, while another might express only Sadness. Annotators were instructed to label all applicable emotions to ensure comprehensive emotional coverage.
3.6.3 Annotation Execution
The annotation of the ANUBHUTI dataset was carried out manually by the same regional dialect experts who were previously involved in the translation of the dataset. Given their linguistic familiarity and contextual awareness, these translators were well-positioned to interpret both the literal meaning and the emotional undertone of the texts within their dialectal context.
Each translator team was responsible for annotating the 500 dialectal sentences corresponding to their assigned region: Mymensingh, Noakhali, Sylhet, and Chittagong. They performed both levels of annotation:
  • Multiclass Annotation: Assigning one thematic category (Political, Religious, or Neutral) per sentence based on its subject matter.
  • Multilabel Annotation: Selecting one or more emotion labels from a predefined set of seven (Anger, Contempt, Disgust, Enjoyment, Fear, Sadness, Surprise), according to the emotional tone conveyed in each sentence.
Table 4: Overview of Dataset Annotation for Chittagong region
Chittagong Target anger contempt disgust enjoyment fear sadness surprise
inU maer/kT Ar sNNGghr/Shr ghTna IyanerO rajoinitk ruup edOyn Ar laI srkarer Aibhnn/dn.ibim/pr entaHr/mii ebyag/Yun Aasaim. Political 1 1 0 0 0 0 0
Engir AaInr shaShn pRitSh/Tha grn ed? Political 0 1 1 0 0 0 0
Aar HNNidn naTk etara girib Neutral 1 0 1 0 0 0 0
Table 4 presents a sample of the final annotation format used in the ANUBHUTI dataset. Each row corresponds to a single sentence translated into the Chittagong dialect and annotated with both a thematic target class (Political, Religious, or Neutral) and multilabel emotional categories. The emotion labels—Anger, Contempt, Disgust, Enjoyment, Fear, Sadness, and Surprise—are represented using binary indicators (1 for presence, 0 for absence), allowing for the expression of multiple emotions per sentence. While the table displays data from only the Chittagong region for illustration, the same annotation schema and format were applied consistently across all four regions: Mymensingh, Noakhali, Sylhet, and Chittagong. This uniform structure ensures compatibility for cross-regional sentiment analysis and enables meaningful comparative modeling.
3.7 Dataset Statistics
The ANUBHUTI dataset comprises a total of 2,000 sentences distributed equally across four major regional dialects of Bangla. As shown in Table 5, each region—Mymensingh, Noakhali, Sylhet, and Chittagong—contributes 500 sentences to ensure a balanced and comparable representation of dialectal variations. This equal allocation supports robust dialect-specific analysis and allows for fair model training and evaluation across regions. The dataset’s regional symmetry was a key design choice to avoid dialectal bias and to ensure methodological consistency in downstream sentiment analysis tasks.
Table 5: Number of Sentences per Region in the ANUBHUTI Dataset
Region Number of Sentences
Mymensingh 500
Noakhali 500
Sylhet 500
Chittagong 500
Total 2000
Figure 2 presents the thematic classification of sentences into three mutually exclusive categories: Political, Religious, and Neutral. This distribution highlights the predominance of political content in the dataset, reflecting the socio-political context prevalent in the source material.
Figure 2: Multiclass Distribution Of Anubhuti Dataset
Figure 3 depicts the frequency of emotion labels assigned in the multilabel annotation scheme. The chart reveals that emotions such as Anger, Contempt, and Disgust occur more frequently, indicating the emotionally charged nature of the regional dialect texts, especially in political and religious discussions. Conversely, less frequent labels like Fear and Surprise reflect more nuanced or less commonly expressed sentiments.
Figure 3: Multilabel Distribution Of Anubhuti Dataset
Together, these figures provide a comprehensive overview of the dataset’s thematic and emotional composition, which is critical for designing and evaluating sentiment analysis models tailored to regional Bangla dialects.
3.8 Data Distribution
The ANUBHUTI dataset was constructed to ensure balanced and representative coverage across multiple dimensions, including regional dialects, emotional sentiment categories, contextual themes, and machine learning data splits. A total of 2,000 dialectal sentences were used in the initial version of the dataset, with 500 sentences allocated per dialect region: Mymensingh, Noakhali, Sylhet, and Chittagong.
Dialectal Distribution
The dataset maintains an equal distribution across the four selected Bangla regional dialects. This facilitates fair comparison in dialect-specific sentiment analysis and ensures that models trained on the dataset are exposed to a balanced variety of regional linguistic patterns.
Sentiment and Contextual Label Distribution
Each sentence was manually annotated using a multilabel sentiment schema comprising seven emotional classes: anger, surprise, contempt, disgust, fear, enjoyment, and sadness. Additionally, a single contextual label was assigned to each sentence in a mutually exclusive fashion: political, religious, or neutral.
This dual-layer annotation allows for comprehensive modeling of both affective and topical dimensions. Care was taken to avoid extreme class imbalance across these categories, thereby supporting reliable supervised learning.
Train-Validation-Test Split
To prepare the dataset for machine learning experiments, the entire annotated dataset was randomly partitioned into three subsets:
  • Training Set (70%): Used to train sentiment classification models.
  • Validation Set (15%): Used to fine-tune model parameters and prevent overfitting.
  • Test Set (15%): Used to evaluate the final model’s generalization performance on unseen data.
Stratified sampling was applied during the splitting process to preserve the proportional distribution of sentiment labels and context classes across all three subsets. This approach ensures that each split contains a representative sample of the emotional and thematic diversity present in the dataset, which is critical for training robust and unbiased models.
4 Limitations
Despite its contributions, the ANUBHUTI dataset has several limitations that should be acknowledged:
  • Limited Domain Coverage: The dataset primarily includes religious and political texts, which may not generalize well to other domains such as sports, entertainment, or daily conversation.
  • Sentiment Class Imbalance: While efforts were made to balance sentiment classes, certain dialects may still exhibit skewed sentiment distributions, which can affect model fairness and generalization.
These limitations open directions for future improvements, such as collecting more naturally occurring dialectal text, expanding the dataset to include diverse domains, and increasing the number of samples per region.
5 Ethics Statements
The data used to construct the ANUBHUTI dataset does not raise any ethical concerns, as it was collected from a publicly available dataset and through Manual Translation. The dataset contains no sensitive or private information, ensuring compliance with ethical standards. Additionally, no data was sourced from personal communications or restricted sources. This dataset is intended solely for NLP research and development, and no human or animal subjects were involved in its creation.
6 Data Availability
The ANUBHUTI dataset, which contains annotated sentiment analysis data in Bangla regional dialects, is publicly available for research and academic purposes. Researchers interested in utilizing the dataset can access it through the following link:Dataset of Sentiment Analysis for Regional Bangla Language (Original data) (Mendeley Data)
7 CRediT Author Statement
Swastika Kundu: Data curation, Writing– original draft, Visualization, Validation
Autoshi Ibrahim: Data curation, Investigation, Writing– original draft
Mithila Rahman: Validation, Software, Conceptualization
Tanvir Ahmed: Supervision, Writing– review & editing
8 Acknowledgements
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
9 Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.